### v1

@@ 6 Conclusion
We introduced TSAP for self-supervised time series anomaly detection, which is the first attempt that automatically (self-)tunes the augmentation hyperparameters on time series data in an unsupervised manner. TSAP includes a differentiable model to augment the input time series data with various anomaly types, and an unsupervised validation loss that assists in aligning the augmented and test data. Experiments showed TSAP’s ability in effectively selecting the augmentation type along with its continuous hyperparameters. Across various real-world datasets with different types of time series anomalies, TSAP outperformed a diverse list of baselines, including modern neural and self-supervised approaches. While being the first self-tuning SSL solution to TSAD, our work opens new research directions, including an expanded catalog of supported anomaly types and possible extension of the ideas to multivariate time series data.
Appendix 0.A Appendix
0.A.1 Types of Time Series Anomalies
We include six types of time series anomalies:
@@ 0.A.1 Types of Time Series Anomalies
Fig. 2 visualizes each type of anomaly and demonstrates how each hyperparameter controls the injection of the anomaly in the time series for platform anomaly.
0.A.2 Dataset Details
PhysioNet
The 2017 PhysioNet Challenge dataset [9] comprises a diverse array of real-world 1-lead 300 Hz ECG recordings. We use ECG recordings, each 9 seconds in length with $K=2700$ time-steps and standardized to have a zero mean and standard deviation of one. Injected anomalies represent 10% of the data. Here, we include three additional controlled TSAD tasks based on the PhysioNet data as shown in Table 3.
Table 4 (top) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in PhysioNet. The hyperparameters location and length are normalized by $K$. Noting that the extremum anomaly always occurs on a single timestamp in the time series, thus length is always $1$.
Table 3: Anomaly profile of additional TSAD tasks E–G.
Profile PhysioNet E PhysioNet F PhysioNet G
Type Mean shift Mean shift Extremum
Level Fixed Fixed Fixed
Location Random Random Random
Length Random Fixed N/A
MoCap
The CMU Motion Capture (MoCap) dataset [^3] http://mocap.cs.cmu.edu/ includes signal data from various sensors on subjects’ bodies as they perform different activities (walking, jumping, or running). As we focus on a univariate setting, only the sensor signal on the left femur is used. Since each data contains the time series data with a short length, we stitch them into a longer one in our experiment. To ensure it is done smoothly, we identified the start and end points of each gait phase for the stitching. We further add random noises to augment the normal samples in the dataset. Each signal is normalized between $-1$ and $1$ and truncated to length $K=1500$. Constructed anomalies represent 10% of the data. As opposed to PhysioNet, we consider MoCap as a dataset with natural or real anomalies. Therefore, we do not create additional TSAD tasks from this dataset.
Table 4 (bottom) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in MoCap. The hyperparameters location and length of platform anomaly are normalized by $K$. The hyperparameters location and length of frequency anomaly denote the starting gait phase and the length of gait phases, respectively.
Table 4: Hyperparameter space for each type of anomaly.
PhysioNet
Anomaly Types location length level
Platform $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
Mean Shift $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
Amplitude $[100..2000]$ $[400..600]$ $\{0.5k+1|k\in[0..10]\}$
Trend $[100..2000]$ $[400..600]$ $\{0.002k-0.01|k\in[0..10]\}$
Extremum/Spike $\{100k|k\in[1..26]\}$$\{1\}$ $\{3k-15|k\in[0..10]\}$
MoCap
Anomaly Types location length level
Platform $[200..800]$ $[100..200]$ $\{0.2k-1|k\in[0..10]\}$
Frequency Shift $[1..3]$ $[1..6]$ $[1..3]$
0.A.3 Model Configurations
Following Sec. 4.1, we provide further details on the configurations for TSAP and all baselines.
TSAP configuration:
In Table 5, we provide a comprehensive overview of the configuration details for the different components of TSAP. Table 5 (top) shows the configuration details for the different components of $f_{\mathrm{aug}}$ following Fig. 1 (left). Similarly, Table 5 (bottom) shows the configuration details for the components of $f_{\mathrm{det}}$ following Fig. 1 (right).
@@ 0.A.3 Model Configurations
Baseline configurations:
We used the parameter settings recommended by the original authors for all baseline models. This approach ensures that our comparisons are fair and align with the intended set-up of these methods originally proposed by the authors.
  • OC-SVM: We use author-recommended hyperparameters based on [17].
  • LOF: We use author-recommended hyperparameters based on [7].
  • IF: We use author-recommended hyperparameters based on [21].
@@
  • SR-CNN: We use author-recommended hyperparameters based on [28].
@@
  • NeuTraL-AD: We use author-recommended hyperparameters based on [26]. We tune augmentation type separately for each dataset using labeled validation data.
0.A.4 Additional Results
@@
We present additional results for continuous and discrete augmentation hyperparameter tuning, as well as two additional ablation studies.
0.A.4.1 A.4.1 Continuous Augmentation Hyperparameter Tuning
@@
In addition to the results on continuous hyperparameter tuning for PhysioNet A and B (see Fig. 3), we demonstrate TSAP’s efficacy in tuning the continuous hyperparameters on five additional TSAD tasks. These include PhysioNet C and D, which feature Trend anomalies (cf. Table 1), as well as PhysioNet E and F, showcasing Mean shift anomalies, and PhysioNet G, illustrating Extremum anomalies (cf. Table 3).
PhysioNet C & D:
@@
Figure 6: (best in color) Tuning the continuous augmentation hyperparameter(s) with TSAP for Trend anomalies. Top: Given Trend anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val. loss (middle), and leading to high detection AUROC performance (right). Bottom: Multiple continuous hyperparameters, here both level and length are accurately tuned to near true values (left), as guided by minimizing the val. loss (middle), achieving high AUROC (right).
@@
Similarly to Trend anomalies, Mean shift anomalies are inherently subtle, especially when the level is close to zero. To this end, we show in Fig. 7 how TSAP properly tunes the continuous hyperparameters for level and length in PhysioNet E and F for several initializations.
PhysioNet G:
@@
Lastly, Fig. 8 showcases TSAP’s ability to tune the level of the spike in the Extremum anomalies while the location is randomized. Note that Extremum anomalies have no length by definition. The ECG recordings in PhysioNet contain many natural spikes. As such, validation loss is low by default. Nonetheless, TSAP successfully tunes two out of four initializations and reflects – though subtly – this difference in the validation loss. Ultimately, this again leads to a well-tuned $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$ that performs strongly on $\mathcal{D}_{\mathrm{val}}$.
@@
Figure 7: (best in color) Tuning the continuous augmentation hyperparameter(s) with TSAP for Mean shift anomalies. Top: Given Mean shift anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val. loss (middle), and leading to high detection AUROC performance (right). Bottom: Multiple continuous hyperparameters, both level and length are accurately tuned to near true values (left), as guided by minimizing the val. loss (middle), achieving high AUROC (right). Mean shift anomalies have low val. loss by default due to the subtle nature of Mean shift anomalies.
@@
Figure 8: (best in color) Tuning the continuous augmentation hyperparameter(s) with TSAP for Extremum anomalies. Given Extremum anomalies at true level (red dashed line), various initializations converge accurately near the true value (left), following the minimized values of val. loss (middle), and leading to high detection AUROC performance (right). Note how Extremum anomalies have low val. loss by default due to the natural presence of spikes in ECG data.
0.A.4.2 A.4.2   Discrete Augmentation Hyperparameter Tuning
@@
We showcased TSAP’s ability to tune the discrete hyperparameter, anomaly type, in Fig. 4 for controlled and natural TSAD tasks. Given the direct applicability and significance of discrete hyperparameter tuning in real-world contexts, we present extended results for natural TSAD tasks by additionally discussing discrete hyperparameter tuning for MoCap B (cf. Table 1).
MoCap B:
@@
Fig. 9 shows the discrete hyperparameter tuning for the unknown anomaly type in MoCap B. TSAP was initialized twice: first with $f_{\mathrm{aug}}$ pre-trained for injecting Frequency shift anomalies, and second with $f_{\mathrm{aug}}$ pre-trained for injecting Platform anomalies. The validation loss (top) indicates a strong alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and the unlabeled $\mathcal{D}_{\mathrm{val}}$ when TSAP is initialized with Frequency shift anomalies. This is also reflected in $f_{\mathrm{det}}$’s performance on $\mathcal{D}_{\mathrm{val}}$ (middle). Visually, we can indeed confirm that Frequency shift anomalies (bottom – red) appear to be more similar to the running pattern (bottom – black) as opposed to Platform anomalies (bottom – purple).
0.A.4.3 A.4.3   Additional Ablation Studies
@@
Figure 9: (best in color) Tuning discrete hyperparameter (anomaly type) with TSAP. For Run anomalies in MoCap B with unknown type (black), val. loss favors Frequency shift (red) that leads to high AUROC (middle) and mimics well true anomaly (bottom), and effectively rejects type Platform (purple) with poor performance.
Following Sec. 4.4, we present additional ablation studies for the second-order optimization within TSAP and the normalization of the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ obtained through $f_{\mathrm{det}}^{\mathrm{enc}}$.
Second-order Opt. Ablation:
@@
Embedding Normalization Ablation:
We show the level-estimation and performance of $f_{\mathrm{det}}$ on PhysioNet C when embedding normalization is disabled in Fig. 10 (bottom). While $a$ initialized at $0.01$ eventually leads to the correct level, the estimation process is much more volatile compared to when normalization is enabled as shown in Fig. 6 (top).
Figure 10: (best in color) Overview of additional ablation studies. Top: We disable second-order optimization in TSAP, leading to a highly unstable estimation process of $a$ (left) and poor performance of $f_{\mathrm{det}}$ (right). Bottom: We disable normalization of the embeddings in TSAP. Estimation of $a$ (left) is volatile and does not converge well, in turn, performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ is poor in most cases (right).

### v2

@@ 4 Conclusion
We introduced TSAP for self-supervised time series anomaly detection, which is the first attempt that automatically (self-)tunes the augmentation hyperparameters on time series data in an unsupervised manner through a differentiable augmentation model and an unsupervised validation loss to help align augmented and test data. While being the first self-tuning SSL solution to TSAD, our work opens avenues for an expanded catalog of anomaly types and extensions to multivariate time series data.
Acknowledgments
Funding: This work was supported by the Research Foundation - Flanders (FWO) [grant number G0C6721N] and [grant number V436123N].
Supplemental Materials
Appendix A Types of Time Series Anomalies
The anomaly generation scheme $g$ accommodates six types of time series anomalies as discussed in Sec. 2.1. Based on $g$, the goal for $f_{\mathrm{aug}}$ is then to learn to inject anomalies into inliers, conditional on the discrete hyperparameter, anomaly type, along with their corresponding hyperparameters as described below. A visual overview of the anomaly types is provided in Fig. 4.
@@ Appendix A Types of Time Series Anomalies
Figure 4: Examples of six different types of time series anomalies; (black) original real-world time series, (red) pseudo anomalies generated by $g$. The goal for $f_{\mathrm{aug}}$ is to learn to inject anomalies into inliers, conditional on the discrete (anomaly type) as well as continuous hyperparameters such as level (magnitude), location (start point) and length (see bottom right).
Appendix B Model Configurations
B.1 Validation Loss
We measure the degree of alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{val}}$ in the embedding space using the Wasserstein distance [10] as described in Sec. 2.2. The Wasserstein distance is a distance measure between probability distributions, of order $p$ between any two marginal distributions $\mu$ and $\nu$, given by:
[…]
B.2 TSAP configuration
The Encoderϕ in $f_{\mathrm{aug}}$ and Encoderθ in $f_{\mathrm{det}}$ are constructed using 1D CNN blocks [12] (transposed 1D CNN for Decoderϕ) for efficient temporal feature extraction. We carefully choose the number of epochs $T$ to allow sufficient time for the convergence of $\mathbf{a}$, with empirical evidence suggesting that $T=100$ typically suffices. For the number of inner-loops $L$, we opt for $L=5$, aligned with [8], such that $f_{\mathrm{det}}$ has adequate time to learn effective discriminative embeddings for $\mathcal{D}_{\mathrm{aug}}$ and $\mathcal{D}_{\mathrm{trn}}$. Table 2 provides a comprehensive overview of the configuration details for the different components of TSAP.
@@ B.2 TSAP configuration
B.3 Baseline configurations
We compare TSAP with a selection of established baselines, including traditional and deep learning-based methods with demonstrated efficacy in TSAD [13]. The traditional methods consist of different modeling approaches; namely, One-Class Support Vector Machines (OC-SVM) [14]; Local Outlier Factor (LOF) [15]; (ARIMA) [16]; Isolation Forest (IF) [17]; and the Matrix Profile (MP) [18]. On the deep learning side, we benchmark against the Encoder-Decoder LSTM (EncDec-LSTM) [19]; the Spectral Residual Convolutional Neural Network (SR-CNN) [20]; the Unsupervised Anomaly Detection (USAD) for TSAD [21]; and a recent time series foundation model (TimeGPT) [22]. Lastly, we include a state-of-the-art competing method which learns augmentations in the embedding space, called Neural Transformation Learning for (TS)AD (NeuTraL-AD) [7]. This diverse set of baselines allows for a comprehensive analysis across different approaches within the TSAD domain. The details for the baseline configurations are provided below. All models were trained on a single NVIDIA Tesla P100 GPU.
    OC-SVM: We use author-recommended hyperparameters [23].
    LOF: We use author-recommended hyperparameters [15].
    IF: We use author-recommended hyperparameters [17].
    ARIMA: We use AutoARIMA to select hyperparameters [16].
    MP: We use author-recommendations to set the window size $m$ [18]
@@ B.3 Baseline configurations
    SR-CNN: We use author-recommended hyperparameters [20].
@@
    NeuTraL-AD: We use author-recommended hyperparameters [7]. We tune augmentation type separately for each dataset using labeled validation data.
    TimeGPT: We tune the confidence interval [22] for each dataset using labeled validation data.
Appendix C Experimental Setup
C.1 Dataset Details
PhysioNet
The 2017 PhysioNet Challenge dataset [9] comprises a diverse array of real-world 1-lead 300 Hz ECG recordings. We use ECG recordings, each 9 seconds in length with $K=2700$ time-steps and standardized to have a zero mean and standard deviation of one. Injected anomalies represent 10% of the data. We include a total of seven controlled TSAD tasks based on the PhysioNet data as shown in Table 3, with PhysioNet A-D discussed in Sec. 3 and the remainder in Appx. D. For example in PhysioNet A and B, given the anomaly type (Platform), the task is to infer or tune, respectively, the hyperparameter(s) level only and both level and length, while anomaly location is random (hence not tuned). For PhysioNet C and D, the respective tuning tasks are the same but for a different anomaly type (Trend). Finally, Table 4 (top) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in PhysioNet. The hyperparameters location and length are normalized by $K$. Noting that the extremum anomaly always occurs on a single timestamp in the time series, thus length is always $1$.
Table 3: Anomaly profile of different TSAD tasks.
Dataset Type Level Location Length
PhysioNet PhysioNet A Platform Fixed Random Random
PhysioNet B Platform Fixed Random Fixed
PhysioNet C Trend Fixed Random Random
PhysioNet D Trend Fixed Random Fixed
PhysioNet E Mean shift Fixed Random Random
PhysioNet F Mean shift Fixed Random Fixed
PhysioNet G Extremum Fixed Random N/A
MoCap MoCap A Jump Fixed Random Random
MoCap B Run Fixed Random Random
MoCap
The CMU Motion Capture (MoCap) dataset [^2] http://mocap.cs.cmu.edu/ includes signal data from various sensors on subjects’ bodies as they perform different activities (walking, jumping, or running). As we focus on a univariate setting, only the sensor signal on the left femur is used. We consider the walking signal as normal data, and the signals of jumping and running as anomalies. To generate normal signals, we stitch the walking signals by identifying the start and end points of each gait phase and add random noise; whereas to generate anomalous ones, we stitch walking or running signals at a random location in the normal signal. We further add random noises to augment the normal samples in the dataset. Each signal is normalized between $-1$ and $1$ and truncated to length $K=1500$. This yields two distinct TSAD tasks as shown in Table 3. For each, constructed anomalies represent 10% of the data. Different from PhysioNet A–G where we only tune the continuous hyperparameter(s) for the given (discrete) anomaly type, for MoCap A and B, we aim to tune both the unknown anomaly type that corresponds to Jump and Run behavior, respectively, as well as the (continuous) hyperparameter level while location and length take random values. Finally, Table 4 (bottom) shows the hyperparameter spaces used to train $f_{\mathrm{aug}}$ in MoCap. The hyperparameters location and length of platform anomaly are normalized by $K$. The hyperparameters location and length of frequency anomaly denote the starting gait phase and the length of gait phases, respectively.
Table 4: Hyperparameter space for each type of anomaly.
PhysioNet
Anomaly Types location length level
Platform $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
Mean Shift $[100..2000]$ $[400..600]$ $\{0.2k-1|k\in[0..10]\}$
Amplitude $[100..2000]$ $[400..600]$ $\{0.5k+1|k\in[0..10]\}$
Trend $[100..2000]$ $[400..600]$ $\{0.002k-0.01|k\in[0..10]\}$
Extremum/Spike $\{100k|k\in[1..26]\}$$\{1\}$ $\{3k-15|k\in[0..10]\}$
MoCap
Anomaly Types location length level
Platform $[200..800]$ $[100..200]$ $\{0.2k-1|k\in[0..10]\}$
Frequency Shift $[1..3]$ $[1..6]$ $[1..3]$
C.2 Evaluation Metrics
Our method calculates anomaly scores on an entire sequence level $\mathbf{x}$, similar to [7]. This is a different set-up compared to novelty detection in time series which typically operates on a point level. Detection on a sequence level can be especially important to spot long-term patterns (e.g. Trend anomalies). As such, we use the $F_{1}$ score and the Area Under the Receiver Operating Characteristic Curve (AUROC) as key performance metrics to quantify detection capability of anomalous sequences. All results are reported on the unseen $\mathcal{D}_{\mathrm{test}}$. We determine the optimal $F_{1}$ score, by enumerating all possible thresholds, given by the anomaly scores for a given segment $\mathbf{x}$. We then compute the corresponding precision and recall for each threshold and select those that yield the highest $F_{1}$. As such, AUROC provides a balanced view whereas $F_{1}$ shows the optimal scenario. Both metrics range from 0 to 1, with higher values indicating superior performance.
Appendix D Additional Results
@@
In this section, we present additional results for continuous (D.1) and discrete (D.2) augmentation hyperparameter tuning, as well as several ablation studies (D.3).
D.1 Continuous Augmentation Hyperparameter Tuning
@@
In addition to the results on continuous hyperparameter tuning for PhysioNet A and B (see Fig. 3), we demonstrate TSAP’s efficacy in tuning the continuous hyperparameters on five additional TSAD tasks. These include PhysioNet C and D, which feature Trend anomalies, as well as PhysioNet E and F, showcasing Mean shift anomalies, and PhysioNet G, containing Extremum anomalies (cf. Table 3).
PhysioNet C & D
@@
Figure 5: Tuning the continuous augmentation hyperparameter(s) with TSAP for Trend anomalies. Top: Given Trend anomalies at true level (red dashed line), various initializations converge near the true value (left), following the minimized values of val. loss (center), and leading to high detection AUROC performance (right). Bottom: Multiple continuous hyperparameters, here both level and length are accurately tuned to near true values (left), as guided by minimizing the val. loss (center), achieving high AUROC (right).
@@
Similarly to Trend anomalies, Mean shift anomalies are inherently subtle, especially when the level is close to zero. We show in Fig. 6 how TSAP properly tunes the continuous hyperparameters for level and length in PhysioNet E and F for several initializations.
PhysioNet G
@@
Lastly, Fig. 7 showcases TSAP’s ability to tune the level of the spike in the Extremum anomalies while the location is randomized. Note that Extremum anomalies have no length by definition. The ECG recordings in PhysioNet contain many natural spikes. As such, validation loss is low by default. Nonetheless, TSAP successfully tunes two out of four initializations and reflects – though subtly – this difference in the validation loss. This again leads to a well-tuned $f_{\mathrm{det}}(\cdot;\boldsymbol{\theta}^{*})$ that performs strongly on $\mathcal{D}_{\mathrm{val}}$.
@@
Figure 6: Tuning the continuous augmentation hyperparameter(s) with TSAP for Mean shift anomalies. Top: Given Mean shift anomalies at true level (red dashed line), various initializations converge near the true value (left), following the minimized values of val. loss (center), and leading to high detection AUROC performance (right). Bottom: Multiple continuous hyperparameters, both level and length are accurately tuned to near true values (left), as guided by minimizing the val. loss (center), achieving high AUROC (right). Mean shift anomalies have low val. loss by default due to the subtle nature of Mean shift anomalies.
@@
Figure 7: Tuning the continuous augmentation hyperparameter(s) with TSAP for Extremum anomalies. Given Extremum anomalies at true level (red dashed line), various initializations converge near the true value (left), following the minimized values of val. loss (center), and leading to high detection AUROC performance (right). Note how Extremum anomalies have low val. loss by default due to the natural presence of spikes in ECG data.
D.2 Discrete Augmentation Hyperparameter Tuning
@@
We showcased TSAP’s ability to tune the discrete hyperparameter, anomaly type, in Fig. 3 for controlled and natural TSAD tasks. Given the direct applicability and significance of discrete hyperparameter tuning in real-world contexts, we present extended results for discrete hyperparameter tuning.
MoCap B
@@
Fig. 8 shows the discrete hyperparameter tuning for the unknown anomaly type in MoCap B. TSAP was initialized twice: first with $f_{\mathrm{aug}}$ pre-trained for injecting Frequency shift anomalies, and second with $f_{\mathrm{aug}}$ pre-trained for injecting Platform anomalies. The validation loss (left) indicates a strong alignment between $\mathcal{D}_{\mathrm{trn}}\cup\mathcal{D}_{\mathrm{aug}}$ and the unlabeled $\mathcal{D}_{\mathrm{val}}$ when TSAP is initialized with Frequency shift anomalies. This is also reflected in $f_{\mathrm{det}}$’s performance on $\mathcal{D}_{\mathrm{val}}$ (center). Visually, we can indeed confirm that Frequency shift anomalies (right – red) appear to be more similar to the running pattern (right – black) as opposed to Platform anomalies (right – purple).
@@
Figure 8: Tuning discrete hyperparameter (anomaly type) with TSAP. For Run anomalies in MoCap B with unknown type (black), val. loss favors Frequency shift (red) that leads to high AUROC (center) and mimics well true anomaly (right), and effectively rejects inferior type Platform (purple).
D.3 Ablation Studies
We present four ablation studies on the controlled PhysioNet data to support various design strategies of TSAP: validation loss, self-tuning, second-order optimization, and embedding normalization.
Validation Loss
Figure 9: Overview of ablation studies. Top: TSAP’s $\mathcal{L}_{\mathrm{val}}$ is replaced by a point-wise val. loss leading to an erroneous estimation of $a$ (left) and poor performance (right). Bottom: TSAP’s self-tuning module is disabled, $a$ is now randomized. Val. loss indicates poor alignment, reflected in poor performance.
Randomization vs. Tuning
In Fig. 9 (bottom), the self-tuning module is disabled for PhysioNet C, where level is instead randomized (along with location, and length). We observe substantially higher $\mathcal{L}_{\mathrm{val}}$, indicating poor alignment. In turn, $f_{\mathrm{det}}$ struggles to detect the unlabeled anomalies in $\mathcal{D}_{\mathrm{val}}$, showing the utility of TSAP’s systematic hyperparameter (self-)tuning over random choice.
Second-order Opt. Ablation
@@ D.3 Ablation Studies
Embedding Normalization Ablation
We show the level-estimation and performance of $f_{\mathrm{det}}$ on PhysioNet C when the normalization of the embeddings $\{\mathcal{Z}_{\mathrm{trn}},\mathcal{Z}_{\mathrm{aug}},\mathcal{Z}_{\mathrm{val}}\}$ obtained through $f_{\mathrm{det}}^{\mathrm{enc}}$ is disabled in Fig. 10 (bottom). While $a$ initialized at $0.01$ eventually leads to the correct level, the estimation process is highly volatile compared to when normalization is enabled as shown in Fig. 5 (top).
Figure 10: Overview of additional ablation studies. Top: We disable second-order optimization in TSAP, leading to a highly unstable estimation process of $a$ (left) and poor performance of $f_{\mathrm{det}}$ (right). Bottom: We disable normalization of the embeddings in TSAP. Estimation of $a$ (left) is volatile and does not converge well, in turn, performance of $f_{\mathrm{det}}$ on $\mathcal{D}_{\mathrm{val}}$ is poor in most cases (right).